Serial Problem Business Solutions LO P1, P2, P3, P4, P5Serial Problem Business Solutions LO P1, P2, P3, P4, P5 After the success of the companys first two months, Santana Rey contDec. 2 Paid $950 cash to Hillside Mall for Business Solutions share of mall advertising costs. 3 Paid $480 cash for minor re
Answer #1

Business Solutions Req 1 Journal entries Date Account 2-Dec Advertising expense Credit Calculation Debit $ 950 Cash $ 950 $ 4

Req 2-a Adjusting entries 31-Dec Computer supplies expense Computer Supplies $ 3,295 =2,545+1,400-650 $ 3,295 $ 465 =1,860/12Business Solutions Req 2-b T Accounts Date Debit Credit Balance Date Debit Credit Balance Date Debit Credit Balance Date DebiBusiness Solutions Account Names Cash Accounts Receivable Computer Supplies Prepaid Insurance Prepaid Rent Office Equipment A44,299 Req 4 Business Solutions Income Statement For three months ended Dec 31, 2018 Computer Services Revenue Expenses DepreReq6 Business Solutions Balance Sheet As of Dec 31, 2018 Assets Current Assets Cash Accounts Receivable Computer Supplies PreReq7 Closing entries 31-Dec Computer Services Revenue Income Summary $ 44,299 $ 44,299 $ 16,190 31-Dec Income Summary DepreciBusiness Solutions Account Names Cash Accounts Receivable Computer Supplies Prepaid Insurance Prepaid Rent Office Equipment A